SIMULATION OF THE SPIN-VIBRONIC STRUCTURE IN THE GROUND ELECTRONIC STATE AND EMISSION SPECTRA INTENSITIES FOR CH_{3}O RADICAL.

VADIM L. STAKHURSKY, XIAOYONG LIU, VLADIMIR A. LOZOVSKY^{a}, \underline{ILIAS SIOUTIS}, C. BRADLEY MOORE^{b}, TERRY A. MILLER, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus, OH 43210.

Methoxy is a radical ubiquitous in the reactions of combustion and oxidation of hydrocarbon compounds in the atmosphere. Because of its relatively small size and degenerate X̃^{2}E ground electronic state, it presents a benchmark system for theoreticians to study the multimode Jahn-Teller (JT) effect in the presence of spin-orbit interaction. We have coupled an effective Hamiltonian, including Jahn-Teller coupling terms of up to 3^{d} order, with a powerful computational engine that allows for quick calculations of spin-vibronic eigenstates and, correspondingly, eigenfunctions from which spectrum intensities can be derived. A graphical user interface, coupled with this computational core, helps to visualize the simulated spectra and compare them with the experimental traces. We simulate and discuss the intensities and spin-vibronic level positions, that we have observed in 4 dispersed fluorescence spectra obtained by pumping the Ã state vibrational levels 3^{5}, 3^{1}4^{1}, 3^{3}5^{1} and 3^{3}6^{1}, and complementary data that have been obtained in SEP experiments of Temps and coworkers^{c}.

aDeceased

bCurrent address: Northwestern University, 633 Clark Street, Evanston, IL 60208-1108

cF. Temps, in Molecular Dynamics and Spectroscopy by Stimulated Emission Pumping, edited by Hai-Lung Dai and R. W. Field, World Scientific, NJ, 1995